a 25n force is composed of two different components at a right angle to eachother. if one component is 15N, what is the other component?
1
Expert's answer
2020-10-12T07:46:51-0400

If two components (say, "F_x = 15N" and unknown "F_y" ) are at right angle to eachother, then they satisfiy the Pythagorean's theorem:


"F^2 = F_x^2 + F_y^2"

where "F = 25N" is the magnitude of the force. Expressing the unknown component, obtain:


"F_y = \sqrt{F^2 - F_x^2}\\
F_y = \sqrt{25^2 - 15^2} = 20N"

Answer. 20 N.